OpenStudy (anonymous):

how would I integrate sin2x/2sinx dx

OpenStudy (campbell_st):

well I would use the trig identity \[\sin(2x) = 2\sin(x)\cos(x) \] so the problem is now \[\frac{2\sin(x)\cos(x)}{2\sin(x)}\] then simplify the fraction... might make it a little easier.

OpenStudy (anonymous):

thank you